OpenStudy (anonymous):

find the intercepts and use them to graph the equation 3x+5y=15 I need help starting this problem

OpenStudy (texaschic101):

to find the x intercept, sub in 0 for y 3x + 5y = 15 3x + 5(0) = 15 3x = 15 x = 5 so the x intercept is (5,0) there are 2 ways to find the y intercept.....if you are just looking for the y intercept and not the slope, just sub in 0 for x 3x + 5y = 15 3(0) + 5y = 15 5y = 15 y = 3 so the y intercept is (0,3) but if you needed the slope, you have to put the equation in y = mx + b form, and the m position is the slope and the b position is the y intercept. 3x + 5y = 15 5y = -3x + 15 y = -3/5x + 3 here, the slope is -3/5 and the y intercept is (0,3) You do not have to do it this way unless you want to find the slope as well

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

Wow thank you so much no you did step by step thank you

OpenStudy (texaschic101):

I try to explain the best I can :)